OpenStudy (anonymous):

I need help explaining the solution of this problem... Jessie just finished the probation period at work and got a $0.50 per hour raise and 4 more hours of work per week. Jessie's friend scoffed, "That's only $2 more per week! Who cares?" Why is Jessie's friend wrong about the change in Jessie's weekly pay? What more do we need to know about Jessie's work situation to calculate the actual change in the pay? How is this related to the product rule in calculus class?

OpenStudy (anonymous):

What I have so far which I am not sure if this is right, x=original hours worked y= original new per hour xy=overall pay new pay=(x+4)(y+0.5) =xy+.5x+4y+2 increase=0.5x+4y+2 Can't figure out the product rule applies